the population of village is encreasing at 5% per annum,The population after 2 yrs will become 4410.what was the population in the begining i.e. before 2 yrs
Answer by stanbon(75887)

the population of a village is encreasing at 5% per annum,The population after 2 yrs will become 4410.what was the population in the begining i.e. before 2 yrs
------------
Let "x" be the population 2 years ago.
---
Equation:
(1.05)^2*x = 4410
1.1025*x = 4410
---
x = 4410/1.1025
---
x = 4000
================
Cheers,
Stan H.
================